# Learning to Solve Integer Linear Programs with Davis-Yin Splitting

### v3

@@ 5.3 Warcraft shortest path prediction
We build upon the code provided as part of the PyEPO package Tang & Khalil (2022). We use the same architecture for $w_{\Theta}(d)$ for BB-net, PertOpt-net, and DYS-net—a truncated ResNet18 He et al. (2016) as first proposed in Vlastelica et al. (2019). The baseline is the same truncated ResNet18 architecture trained to predict $x^{\star}(d)$ directly. We train each network for 50 epochs, except for the baseline, which is trained for 150 epochs. The initial learning rate is $5\times 10^{-4}$ and it is decreased by a factor of 10 after 30 and 40 epochs respectively. PertOpt-net and DYS-net are trained to minimize the $\ell_{2}$ loss (5), while BB-net is trained to minimize the Hamming loss as described in Vlastelica et al. (2019). Given test data $\{(d_{i},w(d_{i}),x^{\star}(d_{i}))\}_{i=1}^{N}$, we follow Tang & Khalil (2022) and evaluate a fully trained model $w_{\Theta}(d)$ using relative regret:
$\hat{r}=\sum_{i=1}^{N}\frac{r(d_{i})}{w(d_{i})^{\top}x^{\star}(d_{i})}$ (28)
@@ Results
The results for this experiment are shown in Table 2. Interestingly, BB-net and PertOpt-net are much more competitive in this experiment than in the experiements of Sections 5.1.1, 5.1.2, and 5.2. We attribute this to the “discrete” nature of the true cost vector $w(d)$—for the Warcraft problem entries of $w(d)$ can only take on four, well-separated values—as opposed to the more “continuous” nature of $w(d)$ in the previous experiments. This difference is illustrated in Figure 5.
@@
Algorithm Test Relative Regret Prediction Accuracy Time (in hours)
Baseline $19.1150$ $0.0\%$ $0.07$
BB-net $0.1435$ $91.43\%$ $5.56$
PertOpt-net $0.4566$ $35.71\%$ $3.15$
DYS-net $0.2745$ $51.43\%$ $4.13$
Table 2: Results for the 12-by-12 Warcraft shortest path prediction problem. The low relative regret of BB-net could be attributed to the “discrete” nature of the true cost vectors as well as its use of the Hamming loss in training.
@@ 6 Conclusions
This work presents a new method for learning to solve ILPs using Davis-Yin splitting which we call DYS-net. We prove that the gradient approximation computed in the backward pass of DYS-net is indeed a descent direction, thus advancing the current understanding of implicit networks. Our experiments show that DYS-net is capable of scaling to truly large ILPs, and outperforms existing state-of-the-art methods in some settings. Our experiments also reveal an interesting dichotomy between problems in which entries of $w(d)$ may take on only a handful of discrete values and problems in which $w(d)$ is more “continuous”. Future work could explore this dichotomy further, as well as apply DYS-net to additional ILPs, for example the traveling salesman problem.
@@ Appendix A Appendix A: Proofs
Theorem 2. Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be as in (14), and suppose $f_{\Theta}(x;\gamma,d)=w_{\Theta}(d)^{\top}x+\frac{\gamma}{2}\|x\|_{2}^{2}$ for any neural network $w_{\Theta}(d)$. For any $\alpha\in(0,2/\gamma)$ define the sequence $\{z^{k}\}$ by:
@@
If $x^{k}\triangleq P_{\mathcal{C}_{2}}(z^{k})$ then $\|x^{k+1}-x^{k}\|^{2}_{2}=O(1/k)$.
@@
Note that (31) can equally be viewed as a variational inequality with operator $F_{\Theta}(x;d)=\nabla_{x}f_{\Theta}({x}_{\Theta};d)$. We deduce that $z_{\Theta}$ with $x_{\Theta}=P_{\mathcal{C}_{2}}(z_{\Theta})$ is a fixed point of
$T_{\Theta}(z)$ $\triangleq z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\left(2\cdot P_{\mathcal{C}_{2}}(z)\!-\!z\!-\!\alpha\left[w_{\Theta}(d)+\gamma P_{\mathcal{C}_{2}}(z)\right])\right)$
$=z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\left(\left(2-\alpha\gamma\right)\cdot P_{\mathcal{C}_{2}}(z)-z-\alpha w_{\Theta}(d)\right)$ (32a)
from (McKenzie et al., 2021, Theorem 4.2), which is itself a standard application of ideas from operator splitting Davis & Yin (2017); Ryu & Yin (2022).
As $\nabla_{x}f_{\Theta}({x}_{\Theta};\gamma,d)$ is $\gamma$-Lipschitz continuous it is also $\nabla_{x}f_{\Theta}$ is $1/\gamma$-cocoercive by the Baillon-Haddad theorem Baillon & Haddad (1977); Bauschke & Combettes (2009). (McKenzie et al., 2021, Theorem 3.3) then implies that $z^{k}\to z_{\Theta}$ but does not give a convergence rate. In fact, the convergence rate can be deduced from known results; because $T_{\Theta}$ is averaged for $\alpha<2/\gamma$ the rate follows from (Ryu & Yin, 2022, Theorem 1). ∎
@@
Differentiating the expression for $T_{\Theta}$ in (30) with respect to $z$ yields
@@
$\left.\frac{\partial T_{\Theta}}{\partial z}\right|_{z=\hat{z}}$ $={\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}\pgfsys@color@rgb@stroke{0}{0.5}{0}\pgfsys@color@rgb@fill{0}{0.5}{0}I-P_{\mathcal{H}_{2,\hat{z}}}}+(2-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}}-P_{\mathcal{H}_{1}}$ (42a)
$={\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}\pgfsys@color@rgb@stroke{0}{0.5}{0}\pgfsys@color@rgb@fill{0}{0.5}{0}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}+(2-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}\left(I-P_{\mathcal{H}_{2,\hat{z}}^{\bot}}\right)-P_{\mathcal{H}_{1}}$ (42b)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}+P_{\mathcal{H}_{1}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}-P_{\mathcal{H}_{1}}$ (42c)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}(I-P_{\mathcal{H}_{1}})P_{\mathcal{H}_{2,\hat{z}}^{\top}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}(I-P_{\mathcal{H}_{2,\hat{z}}^{\bot}})}$ (42d)
@@
We first rewrite $\mathcal{H}_{1}^{\perp}$ and $\mathcal{H}_{2,{z}_{\Theta}}^{\perp}$. The subspace $\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$ is spanned by all non-positive coordinates of ${z}_{\Theta}$. By (17), $[{x}_{\Theta}]_{i}=\max\{0,[{z}_{\Theta}]_{i}\}$, and so $i\in\mathcal{A}({x}_{\Theta})$ if and only if $[{z}_{\Theta}]_{i}\leq 0$. It follows that
@@
noting the final equality holds since $w_{1}$ and $w_{2}$ are orthogonal. Because (A) holds for arbitrarily chosen $w\in{\mathbb{R}}^{n}$,
@@
From the proof of Theorem 6 we see that $T_{\Theta}$ is contractive with constant $\Gamma=\sqrt{\max\{\cos(\tau),(1-\alpha\gamma)^{2}\}}$ and so the main theorem of Fung et al. (2022), guaranteeing $p_{\Theta}$ is a descent direction, as long as the condition number of $(\partial T_{\Theta}/\partial\Theta)^{\top}(\partial T_{\Theta}/\partial\Theta)$ is less than $1/\Gamma$. ∎
@@
Similar guarantees, albeit with less restrictive assumptions on $\partial T_{\Theta}/\partial\Theta$, can be deduced from the results of the recent work Bolte et al. (2024).
@@ Appendix B Derivation for Canonical Form of Knapsack Prediction Problem
For completeness, we explain how to transform the $k$-knapsack prediction problem into the canonical form (8), and how to derive the standardized representation of the constraint polytope $\mathcal{C}$. Recall that the $k$-knapsack prediction problem, as originally stated, is
@@ Appendix C Experimental Details
C.1 Additional Data Details for Knapsack Problem
As mentioned, we use PyEPO Tang & Khalil (2022) to generate the training data. Specifically, $d\in{\mathbb{R}}^{5}$ is sampled from the multivariate Gaussian distribution with mean 00 and covariance $I$. Then, $B\in{\mathbb{R}}^{n\times 5}$ is sampled where each $B_{ij}=+1$ with probability $0.5$ and $-1$ with probability $0.5$. The associated cost vector $w(d)$ is computed as
@@ C.1 Additional Data Details for Knapsack Problem
where $\mathrm{deg}=4$ and $\epsilon_{ij}$ is sampled uniformly from the interval $[0.5,1.5]$.
C.2 Additional Training Details for Knapsack Problem
For all models we use an initial learning rate of $10^{-3}$ and a scheduler that reduces the learning rate whenever the validation loss plateaus. We also used weight decay with a parameter of $5\times 10^{-4}$. All networks were trained using a AMD Threadripper Pro 3955WX: 16 cores, 3.90 GHz, 64 MB cache, PCIe 4.0 CPU and an NVIDIA RTX A6000 GPU.
C.3 Additional Model Details for Large-Scale Shortest Path Problem
Our implementation of PertOpt-net used a PyTorch implementation [^8] See code at github.com/tuero/perturbations-differential-pytorch of the original TensorFlow code [^9] See code at github.com/google-research/google-research/tree/master/perturbations associated to the paper Berthet et al. (2020). We experimented with various hyperparameter settings for 5-by-5 grids and found setting the number of samples equal to 3, the temperature (i.e. $\varepsilon$) to 1 and using Gumbel noise to work best, so we used these values for all other shortest path experiments. Our implementation of BB-net uses the blackbox-backprop package [^10] See code at https://github.com/martius-lab/blackbox-backprop associated to the paper Pogančić et al. (2019). We found setting $\lambda=5$ (see Pogančić et al. (2019) for a description of this hyperparameter) to work best for 5-by-5 grids, so we use this value for all other shortest path experiments.
C.4 Additional Training Details for Shortest Path

### v4

@@ 5.3 Warcraft shortest path prediction
We build upon the code provided as part of the PyEPO package (Tang & Khalil, 2022). We use the same architecture for $w_{\Theta}(d)$ for BB-net, PertOpt-net, and DYS-net—a truncated ResNet18 (He et al., 2016) as first proposed in Vlastelica et al. (2019). We train each network for 50 epochs, except for the baseline. The initial learning rate is $5\times 10^{-4}$ and it is decreased by a factor of 10 after 30 and 40 epochs respectively. PertOpt-net and DYS-net are trained to minimize the $\ell_{2}$ loss equation 5, while BB-net is trained to minimize the Hamming loss as described in Vlastelica et al. (2019). We evaluate the quality of $w_{\Theta}(d)$ using normalized regret equation 25.
@@ Results
The results for this experiment are shown in Table 2. Interestingly, BB-net and PertOpt-net are much more competitive in this experiment than in the experiments of Sections 5.1.1, 5.1.2, and 5.2. We attribute this to the “discrete” nature of the true cost vector $w(d)$—for the Warcraft problem entries of $w(d)$ can only take on four, well-separated values—as opposed to the more “continuous” nature of $w(d)$ in the previous experiments. Thus, an algorithm that learns a rough approximation to the true cost vector will be more successful in the Warcraft problem than in the shortest path problems of Sections 5.1.1 and 5.2. This difference is illustrated in Figure 5. We also note that model selection using a validation set is beneficial for all approaches, but particularly so for BB-net and DYS-net, which achieve their best-performing models in the first several epochs of training.
@@
Algorithm Test Normalized Regret Time (in hours)
BB-net $0.1204$ $0.11$
PertOpt-net $0.1089$ $2.07$
DYS-net $0.0889$ $0.09$
CVX-net $0.0214$ $2.33$
Table 2: Results for the 12-by-12 Warcraft shortest path prediction problem. We select the model achieving best normalized regret on the validation set. The time displayed is the time till this best normalized regret is achieved. All results are averaged over three runs.
@@ 6 Conclusions
This work presents a new method for learning to solve ILPs using Davis-Yin splitting which we call DYS-net. We prove that the gradient approximation computed in the backward pass of DYS-net is indeed a descent direction, thus advancing the current understanding of implicit networks. Our experiments show that DYS-net is capable of scaling to truly large ILPs, and outperforms existing state-of-the-art methods, at least when training data is of the form $(d,x^{\star}(d))$. We note that in principle DYS-net may be applied given training data of the form $(d,w(d))$; a common paradigm for many predict-then-optimize problems Mandi et al. (2023). However, preliminary experiments showed that SPO+ Elmachtoub & Grigas (2022) achieved a substantially lower regret than DYS-net in this setting. We leave the extension of DYS-net to this data type to future work. Our experiments also reveal an interesting dichotomy between problems in which entries of $w(d)$ may take on only a handful of discrete values and problems in which $w(d)$ is more “continuous”. Future work could explore this dichotomy further, as well as apply DYS-net to additional ILPs, for example the traveling salesman problem.
@@ Appendix A Appendix A: Proofs
Theorem 2. Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be as in equation 14, and suppose $f_{\Theta}(x;\gamma,d)=w_{\Theta}(d)^{\top}x+\frac{\gamma}{2}\|x\|_{2}^{2}$ for any neural network $w_{\Theta}(d)$. For any $\alpha\in(0,2/\gamma)$ define the sequence $\{z^{k}\}$ by:
@@
If $x^{k}\triangleq P_{\mathcal{C}_{2}}(z^{k})$ then the sequence $\{x^{k}\}$ converges to $x_{\Theta}$ in equation 8 and $\|x^{k+1}-x^{k}\|^{2}_{2}=O(1/k)$.
@@
Note that equation 30 can equally be viewed as a variational inequality with operator $F_{\Theta}(x;d)=\nabla_{x}f_{\Theta}({x}_{\Theta};d)$. We deduce that $z_{\Theta}$ with $x_{\Theta}=P_{\mathcal{C}_{2}}(z_{\Theta})$ is a fixed point of
$T_{\Theta}(z)$ $\triangleq z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\left(2\cdot P_{\mathcal{C}_{2}}(z)\!-\!z\!-\!\alpha\left[w_{\Theta}(d)+\gamma P_{\mathcal{C}_{2}}(z)\right])\right)$ (31a)
$=z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\left(\left(2-\alpha\gamma\right)\cdot P_{\mathcal{C}_{2}}(z)-z-\alpha w_{\Theta}(d)\right)$ (31b)
from McKenzie et al. (2021, Theorem 4.2), which is itself a standard application of ideas from operator splitting (Davis & Yin, 2017; Ryu & Yin, 2022).
As $\nabla_{x}f_{\Theta}({x}_{\Theta};\gamma,d)$ is $\gamma$-Lipschitz continuous it is also $\nabla_{x}f_{\Theta}$ is $1/\gamma$-cocoercive by the Baillon-Haddad theorem (Baillon & Haddad, 1977; Bauschke & Combettes, 2009). McKenzie et al. (2021, Theorem 3.3) then implies that $z^{k}\to z_{\Theta}$ and $x^{k}\to x_{\Theta}$. However, this theorem does not give a convergence rate. In fact, the convergence rate can be deduced from known results; because $T_{\Theta}$ is averaged for $\alpha<2/\gamma$ the rate $\|z^{k+1}-z^{k}\|=\mathcal{O}(1/k)$ follows from Ryu & Yin (2022, Theorem 1). Thus,
$\|x^{k+1}-x^{k}\|=\|P_{{\mathbb{C}}_{2}}(z^{k+1})-P_{{\mathbb{C}}_{1}}(z^{k})\|\leq\|z^{k+1}-z^{k}\|=\mathcal{O}(1/k),$ (32)
as desired.
@@
Differentiating the expression for $T_{\Theta}$ in equation 16 with respect to $z$ yields
@@
$\left.\frac{\partial T_{\Theta}}{\partial z}\right|_{z=\hat{z}}$ $={\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}I-P_{\mathcal{H}_{2,\hat{z}}}}+(2-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}}-P_{\mathcal{H}_{1}}$ (42a)
$={\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}+(2-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}\left(I-P_{\mathcal{H}_{2,\hat{z}}^{\bot}}\right)-P_{\mathcal{H}_{1}}$ (42b)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}+P_{\mathcal{H}_{1}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}-P_{\mathcal{H}_{1}}$ (42c)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(I-P_{\mathcal{H}_{1}})P_{\mathcal{H}_{2,\hat{z}}^{\top}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}(I-P_{\mathcal{H}_{2,\hat{z}}^{\bot}})}$ (42d)
@@
We first rewrite $\mathcal{H}_{1}^{\perp}$ and $\mathcal{H}_{2,{z}_{\Theta}}^{\perp}$. The subspace $\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$ is spanned by all non-positive coordinates of ${z}_{\Theta}$. By equation 17, $[{x}_{\Theta}]_{i}=\max\{0,[{z}_{\Theta}]_{i}\}$, and so $i\in\mathcal{A}({x}_{\Theta})$ if and only if $[{z}_{\Theta}]_{i}\leq 0$. It follows that
@@
noting the final equality holds since $w_{1}$ and $w_{2}$ are orthogonal. Because (52) holds for arbitrarily chosen $w\in\mathbb{R}^{n}$,
@@
From the proof of Theorem 6 we see that $T_{\Theta}$ is contractive with constant $\Gamma=\sqrt{\max\{\cos(\tau),(1-\alpha\gamma)^{2}\}}$ and so the main theorem of (Fung et al., 2022), guaranteeing $p_{\Theta}$ is a descent direction, as long as the condition number of $(\partial T_{\Theta}/\partial\Theta)^{\top}(\partial T_{\Theta}/\partial\Theta)$ is less than $1/\Gamma$. ∎
@@
Similar guarantees, albeit with less restrictive assumptions on $\partial T_{\Theta}/\partial\Theta$, can be deduced from the results of the recent work (Bolte et al., 2024).
@@ Appendix B Derivation for Canonical Form of Knapsack Prediction Problem
For completeness, we explain how to transform the $k$-knapsack prediction problem into the canonical form equation 8, and how to derive the standardized representation of the constraint polytope $\mathcal{C}$. Recall that the $k$-knapsack prediction problem, as originally stated, is
@@ Appendix C Experimental Details
C.1 Additional Data Details for PyEPO problems
We use PyEPO (Tang & Khalil, 2022), specifically the functions data.shortestpath.genData and data.knapsack.genData to generate the training data. Both functions sample $B\in\mathbb{R}^{n\times 5}$ with $B_{ij}=+1$ with probability $0.5$ and $-1$ with probability $0.5$ and then construct the cost vector $w(d)$ as
@@ C.1 Additional Data Details for PyEPO problems
where $\mathrm{deg}=4$ and $\epsilon_{ij}$ is sampled uniformly from the interval $[0.5,1.5]$. Note that $w(d)$ is, by construction, non-negative.
C.2 Additional Model Details for PyEPO problems
For PertOpt-net and BBOpt-net we use the default hyperparameter values provided by PyEPO, namely $\lambda=5$ for BBOpt-net, number of samples equal 3, $\epsilon=1$ and Gumbel noise for PertOpt-net. See Pogančić et al. (2019) and Berthet et al. (2020) respectively for descriptions of these hyperparameters. We do so as these are selected via hyperparameter tuning Tang & Khalil (2022). As DYS-net and CVX-net solve a regularized form of the underlying LP equation 7, the regularization parameter $\gamma$ needs to be set. After experimenting with different values, we select $\gamma=5\times 10^{-4}$ for DYS-net and $\gamma=5\times 10^{-1}$ for CVX-net.
C.3 Additional Training Details for Knapsack Problem
For all models we use an initial learning rate of $10^{-3}$ and a scheduler that reduces the learning rate whenever the validation loss plateaus. We also used weight decay with a parameter of $5\times 10^{-4}$.
C.4 Additional Model Details for Large-Scale Shortest Path Problem
Our implementation of PertOpt-net used a PyTorch implementation [^6] See code at github.com/tuero/perturbations-differential-pytorch of the original TensorFlow code [^7] See code at github.com/google-research/google-research/tree/master/perturbations associated to the paper Berthet et al. (2020). We experimented with various hyperparameter settings for 5-by-5 grids and found setting the number of samples equal to 3, the temperature (i.e. $\varepsilon$) to 1 and using Gumbel noise to work best, so we used these values for all other shortest path experiments. Our implementation of BB-net uses the blackbox-backprop package [^8] See code at https://github.com/martius-lab/blackbox-backprop associated to the paper Pogančić et al. (2019). We found setting $\lambda=100$ to work best for 5-by-5 grids, so we use this value for all other shortest path experiments. Again, we select $\gamma=5\times 10^{-4}$ for DYS-net and $\gamma=5\times 10^{-1}$ for CVX-net
C.5 Additional Training Details for Large-Scale Shortest Path Problem